OpenStudy (anonymous):

Solve this system of equations. Write your answer as an ordered pair. 5x + 4y = -14 3x + 6y = 6

OpenStudy (bradely):

multiply the first equation by (6) 30x+24y=-84 -----(1) multiply thesecond equation by (4) 12x+24y=24 -----(2) -------- (1)-(2) 18x=-108 x=-6 Substitute in first equation 5(-6)+4y=-14 4y=16 y=4 If any doubts contact me www.mathskey.com

OpenStudy (anonymous):

needs to be an ordered pair

OpenStudy (bradely):

(-6,4)

OpenStudy (bradely):

this is the ordered pair

OpenStudy (anonymous):

x= -6. y= 4.